Find approximately in minute the inclination to the horizon of an incline which rises 7/6metres in 210 metres.

Grade:11

1 Answers

AKASH KUMAR ROY
6 years ago
The perpendicular height is 7/6metres.
The horizontal lenght is 210metres.
tanx=7/6/210
tanx=1/180
x=tan-1(1/180).